There are 2 wires, i_1 = 3A () and i_2 = 5A (). Th

The symple (o) means out of the page
Answer #1

Magnetic field on point P:

B=\frac{u_{0}}{2\pi }(\frac{I_{1}}{r_{1}}+\frac{I_{2}}{r_{2}})

B=\frac{4\pi X10^{-7}}{2\pi }(\frac{3}{0.1}+\frac{5}{\sqrt{0.1^{2}+0.15^{2}}})

\boldsymbol{B=11.55X10^{-6}\: T}

---------------------------

B)

The force between two perpendicular current carrying wires is zero. Only the sides of the square parallel to the inifinite wire contribute to the force

Force on loop by wire 1

F_{1}=h*I*B_{1}

F_{1}=1.3*100*(\frac{(4\pi X10^{-7})*3}{2\pi *(12.5+0.15)}+\frac{(4\pi X10^{-7})*3}{2\pi *(12.5+0.15+0.5)})

F_{1}=12.10X10^{-6}N

By wire 2

F_{2}=1.3*100*(\frac{(4\pi X10^{-7})*5}{2\pi *(12.5)}+\frac{(4\pi X10^{-7})*5}{2\pi *(12.5+0.5)})

F_{2}=20.4X10^{-6}N

Total force is the sum of 1 and 2

\boldsymbol{F_{T}=32.5X10^{-6}N}